OpenStudy (anonymous):

factor the expression and use the fundamental identities to simplify tan^2x - tan^2xsin^x

OpenStudy (anonymous):

Sorry that's sin^2x

myininaya (myininaya):

\[\tan^2x(1-\sin^2x)\]

myininaya (myininaya):

\[\tan^2x(\cos^2x)\]

myininaya (myininaya):

\[\frac{\sin^2x}{\cos^2x}*\cos^2x=\sin^2x\]

OpenStudy (anonymous):

Thank you.

myininaya (myininaya):

np